# Learning to See and Act: Task-Aware View Planning for Robotic Manipulation

### v1

@@ Introduction
Stage 2: In the second stage, we refine the Multi-View Exploration Policy (MVEP), using the fixed-view TAVP model trained in Stage 1 as a baseline. MVEP is optimized via the Proximal Policy Optimization (PPO) algorithm [93], which typically requires interaction with a physical environment to obtain reward feedback. To mitigate the high time cost of such interaction, we introduce a novel pseudo-environment interaction mechanism. The details of the mechanism are presented in the supplementary materials.
Specifically, we treat the fixed-view TAVP as a reference model to supervise the training of MVEP. Given the same observations and instructions, the reference model yields losses $\mathcal{L}_{\text{ref}}=[\mathcal{L}_{\text{hf}},\mathcal{L}_{\text{rot}},\mathcal{L}_{\text{gri}},\mathcal{L}_{\text{col}}]$. In contrast, MVEP generates dynamic viewpoints for action prediction, resulting in losses $\mathcal{L}_{\text{TAVP}}=[\mathcal{L}^{\prime}_{\text{hf}},\mathcal{L}^{\prime}_{\text{rot}},\mathcal{L}^{\prime}_{\text{gri}},\mathcal{L}^{\prime}_{\text{col}}]$. Taking the reference model as a performance lower bound, our objective is to make $\mathcal{L}_{\text{TAVP}}$ smaller than $\mathcal{L}_{\text{ref}}$. The first reward term is therefore defined as:
$r_{0}=\mathcal{L}_{\text{ref}}-\mathcal{L}_{\text{TAVP}}.$ (6)
@@
where $w_{i}$ are learnable component weights and $\mathcal{N}$ is an online normalization operator. The adaptive normalization maintains the mean and variance of each component using the Welford algorithm [121] for normalization, and clips the final reward to the range [10,10][-10,10] to ensure training stability.
In this stage, only MVEP is trainable, and all the other components of TAVP are frozen.
Stage 3: In the third stage, to enable MVEP to better adapt to action generation, we further fine-tune the entire TAVP model, excluding MVEP, using the same loss functions as in Stage 1.
Method Avg. Success Close Jar Drag Stick Insert Peg Meat off Grill Open Drawer Place Cups Place Wine Push Buttons
RVT2 81.4 100.0 99.0 40.0 99.0 74.0 38.0 95.0 100.0
ARP 81.6 97.6 88.0 53.2 96.0 90.4 48 92.0 100.0
ARP+ 84.9 95.2 99.2 78.4 97.6 92.8 48.8 96.0 100.0
TAVP(Ours) 86.7 100.0 100.0 96.0 100.0 88.0 44.0 96.0 100.0
@@
RVT2 66.0 96.0 96.0 88.0 92.0 35.0 80.0 69.0 100.0 99.0
ARP 68.0 99.2 94.4 85.6 98.4 35.2 55.2 76.8 90.4 100.0
ARP+ 69.6 98.4 86.4 89.6 92.8 46.4 63.2 80.0 97.6 96.0
TAVP(Ours) 76.0 100.0 92.0 88.0 100.0 64.0 68.0 60.0 92.0 96.0
Table 1: Performance comparison across 18 tasks in RLBench. We compare the success rate of each individual task and the average success rate across all tasks between dense fixed-view baselines such as RVT2, ARP, and ARP+ and our TAVP.
Experiments
To fully evaluate our Task-aware View Planning (TAVP) model, we conducted comprehensive experiments in both simulated and real-world environments. Detailed experimental settings for the real-world and the corresponding results are presented in the supplementary materials.
Experimental Setup
Simulation Setup.
@@
Our simulation experiments are conducted on RLBench dataset [19], a multi-task learning benchmark implemented in the CoppeliaSim simulator featuring a 7-degree-of-freedom (7-DoF) Franka Emika Panda robot in a tabletop environment. We employ a curated subset comprising 18 distinct manipulation tasks, each containing multiple variations (tasks details can be found in the supplementary materials). Observations consist of RGB-D images ($128\times 128$) captured from four fixed views: frontal, left shoulder, right shoulder, and wrist-mounted. The action space represents the end-effector pose as $\mathbf{a}=(\mathbf{p}_{\text{xyz}},\mathbf{q}_{\text{rot}},g_{\text{gripper}},c_{\text{collision}})\in\mathbb{R}^{7}\times\{0,1\}^{2}$, where $\mathbf{p}_{\text{xyz}}\in\mathbb{R}^{3}$ denotes Cartesian coordinates, $\mathbf{q}_{\text{rot}}\in\mathbb{R}^{4}$ is the quaternion rotation, $g_{\text{gripper}}\in\{0,1\}$ controls gripper state, and $c_{\text{collision}}\in\{0,1\}$ indicates collision.
Implementation Settings.
@@
In our experiments, the number of camera viewpoints $K=3$, the number of gates in TaskMoE $N_{G}=8$, and the number of task experts $N_{E}=16$. For each task, the Top2Top-2 task experts are selected. The default radial constraint for the camera is $r_{\text{min}}=0.75$ m and $r_{\text{max}}=1.3$ m. Our TAVP model was trained using $4\times$ NVIDIA RTX A800 GPUs with 80GB memory. During testing, we sequentially validated each task using $1\times$ NVIDIA RTX A800 GPU and computed the average success rate.
Main Results
Results Across 18 RLBench Tasks. Compared to dense models that employ fixed camera viewpoints for re-rendering, our sparse Task-Aware View Planning (TAVP) model demonstrates significant performance gains in multi-task scenarios. We conduct benchmark evaluations on 18 distinct RLBench tasks, encompassing a diverse range of manipulation challenges. As shown in Table  1, RVT2, ARP, and ARP+ (where ARP and ARP+ differ only in model capacity) are all dense models utilizing fixed viewpoints for re-rendering. In contrast, our model employs the sparse TAVP approach. The results indicate that our model achieves substantially higher success rates across all tasks compared to the strongest dense baseline, ARP+.
Moreover, the per-task analysis reveals that TAVP consistently outperforms dense baselines in the tasks with the susceptibility to occlusion or fine-grained perceptual demands. In these tasks (e.g., put in cupboard and insert peg), TAVP achieves up to a +56% improvement by repositioning cameras to enhance the joint visibility of the target and end-effector. It indicates that the robot’s gripper or neighbouring objects routinely block the main line of sight, and dynamic views are essential for success. Even on seemingly less challenging operations (close jar, meat off grill, and slide block), TAVP registers increments of 4%, indicating that its benefits are not confined to a handful of corner cases. Averaged over all tasks, the success rate rises from 84.9% (ARP+) to 86.7%, confirming that the combination of active view planning and task-aware feature routing delivers domain-wide gains rather than isolated spikes in performance.
Configuration Average Success Rate (%)
TAVP 86.67
w/o TaskMoE 85.56
w/o Active Exploration Fine-tuning
(Random Viewpoints)
8.89
Table 2: Ablation study results in RLBench tasks.
$K$ $r$ PID $T$ PGC PMS CJ SR

### v4

@@ 3.4 Training Strategy
Stage 2: In the second stage, we refine the Multi-View Exploration Policy (MVEP), using the fixed-view TVVE model trained in Stage 1 as a baseline. MVEP is optimized via the Proximal Policy Optimization (PPO) algorithm [schulman2017proximal], which typically requires interaction with a physical environment to obtain reward feedback. To mitigate the high time cost, we introduce a novel pseudo-environment interaction mechanism. The details of the mechanism are presented in the supplementary materials.
Specifically, we treat the fixed-view TVVE as a reference model to supervise the training of MVEP. Given the same observations and instructions, the reference model yields losses $\mathcal{L}_{\text{ref}}=[\mathcal{L}_{\text{hf}},\mathcal{L}_{\text{rot}},\mathcal{L}_{\text{gri}},\mathcal{L}_{\text{col}}]$. In contrast, MVEP generates dynamic viewpoints for action prediction, resulting in losses $\mathcal{L}_{\text{TVVE}}=[\mathcal{L}^{\prime}_{\text{hf}},\mathcal{L}^{\prime}_{\text{rot}},\mathcal{L}^{\prime}_{\text{gri}},\mathcal{L}^{\prime}_{\text{col}}]$. Taking the reference model as a performance lower bound, our objective is to make $\mathcal{L}_{\text{TVVE}}$ smaller than $\mathcal{L}_{\text{ref}}$. The first reward term is therefore defined as:
$r_{0}=\mathcal{L}_{\text{ref}}-\mathcal{L}_{\text{TVVE}}.$ (6)
@@
where $w_{i}$ are learnable component weights and $\mathcal{N}$ is an online normalization operator. The adaptive normalization maintains the mean and variance of each component using the Welford algorithm [welford1962note] for normalization, and clips the reward to the range [10,10][-10,10] to ensure training stability.
In this stage, only MVEP is trainable, and all the other components of TVVE are frozen.
Stage 3: To enable MVEP to better adapt to action generation, we further fine-tune the entire TVVE model, excluding MVEP, using the same loss functions as in Stage 1.
Method Avg. SR (%) $\uparrow$ Avg. Rank $\downarrow$ Close Jar Drag Stick Insert Peg Meat off Grill Open Drawer Place Cups Place Wine Push Buttons
C2F-ARM-BC [james2022coarse] 20.1 9.67 24.0 24.0 4.0 20.0 20.0 0.0 8.0 72.0
PerAct [shridhar2023perceiver] 49.4 7.06 $55.2{\pm}4.7$ $89.6{\pm}4.1$ $5.6{\pm}4.1$ $70.4{\pm}2.0$ $88.0{\pm}5.7$ $2.4{\pm}3.2$ $44.8{\pm}7.8$ $92.8{\pm}3.0$
HiveFormer [guhur2023instruction] 45.0 8.22 52.0 76.0 0.0 80.0 52.0 0.0 80.0 84.0
PolarNet [chen2023polarnet] 46.0 7.06 36.0 92.0 4.0 100.0 84.0 0.0 40.0 96.0
RVT [goyal2023rvt] 62.9 5.28 $52.0{\pm}2.5$ $99.2{\pm}1.6$ $11.2{\pm}3.0$ $88.0{\pm}2.5$ $71.2{\pm}6.9$ $4.0{\pm}2.5$ $91.0{\pm}5.2$ $\mathbf{100.0{\pm}0.0}$
Act3D [gervet2023act3d] 63.2 5.56 $96.8{\pm}3.0$ $80.8{\pm}6.4$ $24.0{\pm}8.4$ $95.2{\pm}1.6$ $78.4{\pm}11.2$ $3.2{\pm}3.0$ $59.2{\pm}9.3$ $93.6{\pm}2.0$
3D Diffuser Actor [ke20253d] 81.3 3.0 $96.0{\pm}2.5$ $\mathbf{100.0{\pm}0.0}$ $65.6{\pm}4.1$ $96.8{\pm}1.6$ $89.6{\pm}4.1$ $24.0{\pm}7.6$ $93.6{\pm}4.8$ $98.4{\pm}2.0$
RVT2 [goyal2024rvt2] 81.4 2.89 $\mathbf{100.0{\pm}0.0}$ $99.0{\pm}1.7$ $40.0{\pm}0.0$ $\mathbf{99.0{\pm}1.0}$ $74.0{\pm}11.8$ $38.0{\pm}4.5$ $\mathbf{95.0{\pm}3.3}$ $\mathbf{100.0{\pm}0.0}$
ARP [zhang2025autoregressive] 81.6 2.83 97.6 88.0 53.2 $96.0$ $\mathbf{90.4}$ $48.0$ $92.0$ $\mathbf{100.0}$
TVVE(Ours) 86.6 2.17 $\mathbf{100.0{\pm}0.0}$ $\mathbf{100.0{\pm}0.0}$ $\mathbf{98.0{\pm}2.8}$ $94.0{\pm}2.8$ $90.0{\pm}2.8$ $\mathbf{54.0{\pm}2.8}$ $92.0{\pm}5.7$ $\mathbf{100.0{\pm}0.0}$
@@
C2F-ARM-BC [james2022coarse] 0.0 4.0 12.0 8.0 16.0 8.0 0.0 0.0 0.0 68.0
PerAct [shridhar2023perceiver] $28.0{\pm}4.4$ $51.2{\pm}4.7$ $84.0{\pm}3.6$ $17.6{\pm}2.0$ $74.0{\pm}13.0$ $16.8{\pm}4.7$ $26.4{\pm}3.2$ $2.4{\pm}2.0$ $52.0{\pm}0.0$ $88.0{\pm}4.4$
HiveFormer [guhur2023instruction] 32.0 68.0 76.0 8.0 64.0 12.0 4.0 0.0 28.0 80.0
PolarNet [chen2023polarnet] 12.0 32.0 84.0 44.0 56.0 12.0 8.0 8.0 52.0 80.0
RVT [goyal2023rvt] $49.6{\pm}3.2$ $88.0{\pm}5.7$ $91.2{\pm}3.0$ $48.0{\pm}5.7$ $81.6{\pm}5.4$ $36.0{\pm}2.5$ $28.8{\pm}3.9$ $26.4{\pm}8.2$ $72.0{\pm}0.0$ $93.6{\pm}4.1$
Act3D [gervet2023act3d] $67.2{\pm}3.0$ $91.2{\pm}6.9$ $95.2{\pm}4.0$ $32.8{\pm}6.9$ $96.0{\pm}2.5$ $29.6{\pm}3.2$ $4.0{\pm}3.6$ $9.6{\pm}6.0$ $86.4{\pm}6.5$ $94.4{\pm}2.0$
3D Diffuser Actor [ke20253d] $\mathbf{85.6{\pm}4.1}$ $96.0{\pm}3.6$ $\mathbf{97.6{\pm}2.0}$ $82.4{\pm}2.0$ $97.6{\pm}3.2$ $44.0{\pm}4.4$ $68.3{\pm}3.3$ $47.2{\pm}8.5$ $84.0{\pm}4.4$ $99.2{\pm}1.6$
RVT2 [goyal2024rvt2] $66.0{\pm}4.5$ $96.0{\pm}0.0$ $96.0{\pm}2.8$ $\mathbf{88.0{\pm}4.9}$ $92.0{\pm}2.8$ $35.0{\pm}2.8$ $\mathbf{80.0{\pm}2.8}$ $\mathbf{69.0{\pm}5.9}$ $\mathbf{100.0{\pm}0.0}$ $99.0{\pm}1.7$
ARP [zhang2025autoregressive] 68.0 $99.2$ 94.4 $85.6$ $98.4$ $35.2$ 55.2 $76.8$ $90.4$ $\mathbf{100.0}$
TVVE(Ours) $74.0{\pm}8.5$ $\mathbf{100.0{\pm}0.0}$ $78.0{\pm}2.8$ $86.0{\pm}2.8$ $\mathbf{100.0{\pm}0.0}$ $\mathbf{62.0{\pm}8.5}$ $74.0{\pm}2.8$ $64.0{\pm}5.7$ $92.0{\pm}5.7$ $\mathbf{100.0{\pm}0.0}$
Table 1: Performance comparison across 18 tasks in RLBench under the multi-view setup. Our method achieves a substantial improvement in the average success rate. Moreover, the lower average rank of 2.17 demonstrates the superior overall performance and better generalization ability of TVVE across all 18 tasks.
Avg. SR (%) $\uparrow$ Avg. Rank $\downarrow$ close jar open drawer sweep to dustpan meat off grill turn tap slide block put in drawer drag stick push buttons stack blocks
GNFactor [ze2023gnfactor] 31.7 3.9 25.3 76.0 28.0 57.3 50.7 20.0 0.0 37.3 18.7 4.0
Act3D [gervet2023act3d] 65.3 2.7 $52.0_{\pm 5.7}$ $84.0_{\pm 8.6}$ $80.0_{\pm 9.8}$ $66.7_{\pm 1.9}$ $64.0_{\pm 5.7}$ $\mathbf{100.0_{\pm 0.0}}$ $54.7_{\pm 3.8}$ $86.7_{\pm 1.9}$ $64.0_{\pm 1.9}$ $0.0_{\pm 0.0}$
3D Diffuser Actor [ke20253d] 78.4 1.7 82.7±1.9 $\mathbf{89.3_{\pm 7.5}}$ $\mathbf{94.7_{\pm 1.9}}$ 88.0±5.7 80.0±8.6 92.0±0.0 77.3±3.8 $\mathbf{98.7_{\pm 1.9}}$ 69.3±5.0 12.0±3.7
TVVE(Ours) $\mathbf{83.2_{\pm 1.7}}$ 1.6 $\mathbf{100.0_{\pm 0.0}}$ $84.0_{\pm 5.7}$ $94.0_{\pm 2.8}$ $\mathbf{92.0_{\pm 0.0}}$ $\mathbf{92.0_{\pm 5.7}}$ $90.0_{\pm 2.8}$ $\mathbf{98.0_{\pm 2.8}}$ $50.0_{\pm 2.8}$ $\mathbf{82.0_{\pm 2.8}}$ $\mathbf{50.0_{\pm 8.5}}$
Table 2: Performance comparison across 10 tasks in RLBench under the single-view setup. Our TVVE achieves a substantial improvement in the average success rate compared to previous methods.
4 Experiments
To comprehensively assess the effectiveness and generalization of our Task-aware Virtual View Exploration(TVVE), we perform extensive experiments across both simulated and real-world environments.
4.1 Datasets and Experimental Setups
Simulation. Our experiments are conducted on RLBench dataset [james2020rlbench] and RLBench-OG.
@@ 4.1 Datasets and Experimental Setups
RLBench is a multi-task learning benchmark implemented in the CoppeliaSim simulator featuring a 7-degree-of-freedom (7-DoF) Franka Emika Panda robot in a tabletop environment. Its observations consist of RGB-D images ($128\times 128$) captured from four fixed views, including frontal, left shoulder, right shoulder, and wrist-mounted. The action space represents the end-effector pose as $\mathbf{a}=(\mathbf{p}_{\text{xyz}},\mathbf{q}_{\text{rot}},g_{\text{gripper}},c_{\text{collision}})\in\mathbb{R}^{7}\times\{0,1\}^{2}$, where $\mathbf{p}_{\text{xyz}}\in\mathbb{R}^{3}$ denotes Cartesian coordinates, $\mathbf{q}_{\text{rot}}\in\mathbb{R}^{4}$ is the quaternion rotation, $g_{\text{gripper}}\in\{0,1\}$ controls gripper state, and $c_{\text{collision}}\in\{0,1\}$ indicates collision. Following previous works [goyal2024rvt2, ze2023gnfactor], we evaluate our method under two standard configurations: (1) a multi-view setup consisting of 18 manipulation tasks (2–60 variations each) with all four camera views, and (2) a single-view setup including 10 tasks observed only from the front RGB-D camera. The specific tasks are provided in the supplementary materials.
RLBench-OG is derived from RLBench and used for validate the robustness of our TVVE under occluded scenarios and its generalization capability in complex environments. RLBench-OG consists of two distinct suites: the Occlusion Suite and the Generalization Suite. We created variants across 10 tasks for evaluation. In the Occlusion Suite, we designed two experimental configurations: 1) Models are both trained and tested directly under the occluded task configurations; 2) Models are trained in the original RLBench task settings and then evaluated in a zero-shot manner under occluded conditions. For the Generalization Suite, models are trained in the original RLBench environment and subsequently evaluated in a zero-shot manner across various generalization settings. Each task is configured with 50 episodes for training. The mean success rate, evaluated on 25 episodes per task, is reported as mean $\pm$ standard deviation across three independent experimental runs. For simulating occlusion, we introduce occluders or rotate manipulated objects to create occlusions from a frontal viewpoint. For generalization, we introduce perturbations by: 1) altering scene lighting, 2) changing the color and texture of the tabletop, 3) modifying the color and texture of the background, 4) adding distractors, and 5) adjusting the pose of the observation camera. For further details on RLBench-OG, please refer to the supplementary material.
Real World. To evaluate the generalization of our TVVE across different robotic platforms, we conducted experiments on both the Franka Research 3 and the Dobot Nova 2. For the Franka Research 3, a single third-person camera (ORBBEC Femto Bolt) was mounted directly in front of the arm. For the Dobot Nova 2, three Intel RealSense depth cameras were used: an overhead lateral camera (D435i), a front lateral camera (D455), and a wrist-mounted camera on the right arm (D405), as illustrated in the supplementary materials. We designed five distinct manipulation tasks: Pick Grape, Stack Bowls, Push Buttons, Collect Fruits, and Put Item In Drawer. Expert demonstrations were collected with 50 episodes per task for training. To improve efficiency, only key frames from the demonstrations were retained. Additional details of the real-world experiments are provided in the supplementary materials.
@@
Implementation Settings. In our experiments, we set the number of camera viewpoints to $K=3$, the number of gates in TaskMoE to $N_{G}=8$, and the number of task experts to $N_{E}=16$. For each task, the top-2 experts are selected. TVVE was trained on $4$ NVIDIA RTX A800 GPUs. During testing, tasks were sequentially evaluated on a single NVIDIA RTX A800 GPU, and the average success rate was computed.
Figure 4: Visualization of our TVVE in the simulation RLBench Environment and Diffusion Policy in the Real-world Environment. (a) In RLBench, we visualize the re-rendering results of dynamic multi-view in the scenes for the Close Jar and Insert Peg tasks, where EEF in the figures denotes the end-effector. Visualizations of additional tasks are provided in the supplementary material. (b) Sample of Inser peg. (c) Sample of Pick Grape in the real world.
Models Avg. SR (%) $\uparrow$ Avg. Rank $\downarrow$ Occlusion 1 Occlusion 2 Light Color Table Color Table Texture Distractor Background Texture Camera Pose
Diffusion Policy [chi2023diffusion] $23.8\pm 1.9$ 3.0 $27.4\pm 0.8$ $3.4\pm 1.1 22$$9\pm 0.8 22.$$5\pm 0.7 22.$$2.6\pm 0.5$ $24.4\pm 1.4$ $4.9\pm 1.2 22$$2.2\pm 1.2$
ARP [zhang2025autoregressive] $63.7\pm 6.1$ 1.8 $73.0\pm 1.6$ $52.6\pm 1.0$ $59.8\pm 0.8$ $62.7\pm 0.5$ $61.3\pm 1.0$ $62.4\pm 1.4$ $68.1\pm 1.6$ $69.7\pm 1.4$
TVVE(Ours) $67.0\pm 6.2$ 1.1 $75.0\pm 1.2$ $58.0\pm 1.1$ $63.7\pm 1.1$ $4.6\pm 0.7 66$$6.8\pm 0.9$ $60.2\pm 0.6$ $74.3\pm 0.9$ $73.2\pm 0.9$
Table 3: Comprehensive evaluation of model performance under occlusion and generalization settings in RLBench-OG, covering both zero-shot and trained scenarios. Detailed per-task success rates are provided in the supplementary material.
4.2 Comparison with State-of-the-art Methods
To comprehensively evaluate the effectiveness of our  TVVE, we conduct comparisons with a broad range of state-of-the-art approaches, including Diffusion Policy [chi2023diffusion], C2F-ARM-BC [james2022coarse], PerAct [shridhar2023perceiver], Hiveformer [guhur2023instruction], GNFactor [ze2023gnfactor], RVT [goyal2023rvt], RVT2 [goyal2024rvt2], Act3D [gervet2023act3d], 3D Diffuser Actor [ke20253d], and ARP [zhang2025autoregressive]. We evaluate them both quantitatively and qualitatively across three benchmarks: RLBench, RLBench-OG, and real-world manipulation tasks.
Comparisons on RLBench. Tables 1 and 2 present the quantitative results on RLBench [james2020rlbench] under the multi-view and single-view setups, respectively. Our proposed TVVE consistently achieves significant improvements in average success rate across both settings. Specifically, TVVE attains an average success rate of 86.6% under the multi-view setup, surpassing the previous state-of-the-art method ARP [zhang2025autoregressive] by approximately 5%. Notably, substantial gains are observed in tasks such as Insert Peg (from $65.6\%$ to $98.0\%$) and Sort Shape (from $36.0\%$ to $62.0\%$). A qualitative example of Insert Peg is illustrated in Fig. 4(a)(b). It shows that compared to ARP [zhang2025autoregressive], which relies on fixed viewpoints, our TVVE dynamically renders more informative perspectives, effectively mitigating occlusions and collisions while providing richer context for more accurate action prediction, validating that dynamic “seeing” underpins robust “acting.” Moreover, as indicated by the lower average rank of 2.17 in Table 1, TVVE demonstrates superior overall performance across all 18 tasks, highlighting its enhanced generalization capability, which is primarily attributed to the proposed TaskMoE design.
Arm Method Avg. SR (%) $\uparrow$ Pick Grape Stack Bowls Push Buttons Collect Fruits Put Item In Drawer
Dobot Diffusion Policy 68.0 90.0 70.0 70.0 50.0 60.0
TVVE (Ours) 88.0 100.0 90.0 100.0 70.0 80.0
Franka Diffusion Policy 58.0 90.0 50.0 60.0 40.0 50.0
ARP 72.0 100.0 70.0 80.0 50.0 60.0
TVVE (Ours) 78.0 100.0 70.0 90.0 60.0 70.0
Table 4: Comparison of task success rates for Dobot and Franka robots across different manipulation tasks. For more real-robot experiments, including robustness and generalization testing, please refer to the supplementary materials. *Franka and Dobot adopt different experimental configurations.
Comparisons on RLBench-OG. Table 3 highlights that TVVE achieves the best overall success rate (67.0% $\pm$ 6.2) and the best average rank (1.1) on RLBench-OG, surpassing ARP (63.7% $\pm$ 6.1, rank 1.8) and the Diffusion Policy baseline (23.8% $\pm$ 1.9, rank 3.0). The advantage of our virtual-view exploration, dynamic virtual multi-view re-rendering and TaskMoE manifests consistently across perturbations: TVVE yields +2.0/+5.4 absolute gains over ARP for Occlusion 1/2, maintains higher resilience to background texture shifts (74.3% $\pm$ 0.9 vs. 68.1% $\pm$ 1.6), and preserves reliable execution under camera pose changes (73.2% $\pm$ 0.9 vs. 69.7% $\pm$ 1.4). These improvements confirm that learning task-aware viewpoints and decoupled visual representations mitigates the interference issues, leading to more robust manipulation under severe OOD disturbances.
Comparisons on Real-World Tasks. In Dobot robot experiment, compared to the baseline method Diffusion Policy (DP) [chi2023diffusion], TVVE achieves superior performance across all tasks, attaining a significantly higher average success rate of 88.0% versus 68.0% for DP. Notable gains are observed in tasks such as “Push Buttons” (+30%) and “Stack Bowls” (+20%), underscoring the effectiveness of integrating task-aware visual planning into the policy framework. A qualitative example of Pick Grape is illustrated in Fig. 4(c). In Franka robot experiment, TVVE achieves the highest average success rate of 78.0%, outperforming both Diffusion Policy (58.0%) and ARP (72.0%). These results further demonstrate its strong adaptability in real-world scenarios. We report the success rates based on 10 trials for each task, with detailed results provided in Table 4. More real-world results are shown in the supplementary material.
  Configuration   Avg. SR (%) $\uparrow$
  TVVE (TaskMoE++MVEP)   86.6
  w/o TaskMoE   85.6
     Fixed Viewpoints   83.3
     Random Viewpoints   8.9
Table 5: Ablation study results on RLBench multi-view setup. Removing TaskMoE or replacing the explored viewpoints attained by MVEP with random or fixed ones degrades manipulation performance compared to the full TVVE with both components.
4.2.1 Ablation Studies.
In this section, we further analyze the impact of the TaskMoE, Multi-Viewpoint Exploration Policy (MVEP), and the hyperparameters $K$ and $r$, supported by quantitative results in Tables 56, and 7. Specifically, the results in Table 5 are obtained from experiments on all 18 RLBench tasks, while those in Tables 6 and 7 are based on five representative tasks selected from RLBench, namely Put Item in Drawer, Turn Tap, Put Groceries in Cupboard, Put Money in Safe, and Close Jar.
Effectiveness of TaskMoE. The TaskMoE module is designed to enhance multi-task generalization. From Table 5, removing TaskMoE from both the feature extraction and action generation components (w/o TaskMoE) leads to a moderate performance drop to 85.6%, underscoring its importance for multi-task representation learning. To further assess its generalization ability to OOD tasks unseen during training, additional results are presented in Table 6. The results show that TaskMoE not only improves performance on trained tasks but also facilitates transfer to unseen ones. Specifically, on tasks encountered during training, the model equipped with TaskMoE achieves an average success rate of 80.8%, outperforming the variant without TaskMoE (72.0%). This demonstrates its ability to learn and execute diverse, known tasks more effectively. For unseen tasks, the model without TaskMoE achieves a success rate of 60.0% on the new task Open Drawer, while the one with TaskMoE attains a higher success rate of 72.0%, highlighting its enhanced generalization capability.
Effectiveness of MVEP. The Multi-Viewpoint Exploration Policy (MVEP) is designed to dynamically explore optimal camera viewpoints that reduce occlusion and provide more informative visual feature for improved robotic manipulation. To evaluate the effectiveness of MVEP, we replace the explored viewpoints with either fixed viewpoints used in RVT2 [goyal2024rvt2] or randomly selected ones. As shown in Table 5, the manipulation success rate drops from $86.6\%$ to $83.3\%$ when using fixed viewpoints, and further plummets to $8.9\%$ when using random viewpoints. This significant degradation demonstrates that the learned viewpoint exploration policy in MVEP effectively identifies informative perspectives, which are crucial for accurate perception and robust action execution.
Seen Unseen
TaskMoE Avg. SR (%) $\uparrow$ PID $T$ PGC PMS CJ
Open drawer
80.8 100.0 100.0 32.0 88.0 84.0 72.0
$\times$ 72.0 96.0 96.0 20.0 64.0 84.0 60.0
Table 6: Experiments with and without TaskMoE. TaskMoE not only improves performance on trained tasks but also enhances generalization to unseen tasks. Here, PID, $T$, PGC, PMS, and CJ denote Put Item in Drawer, Turn Tap, Put Groceries in Cupboard, Put Money in Safe, and Close Jar, respectively.
$K$ $r$ PID $T$ PGC PMS CJ Avg. SR (%) $\uparrow$